I WRITE DOWN THREE CONSECUTIVE EVEN NUMBERS. SIX TIMES THE SMALLEST PLUS TWO TIMES THE LARGEST EQUALS SEVEN TIMES THE MIDDLE NUMBER. FIND THE THREE NUMBERS.
Answer by ankor@dixie-net.com(22740)

I WRITE DOWN THREE CONSECUTIVE EVEN NUMBERS.
x, (x+2), (x+4)
;
SIX TIMES THE SMALLEST PLUS TWO TIMES THE LARGEST EQUALS SEVEN TIMES THE MIDDLE NUMBER.
6x + 2(x+4) = 7(x+2)
6x + 2x + 8 = 7x + 14
8x - 7x = 14 - 8
x = 6 is the smallest number
:
FIND THE THREE NUMBERS
You can do that, I am sure.